On average a University Book Shop sells 300 text books per day with a standard deviation of 50 text books. Find the probability that the University Book Shop will sell at least 400 text books per day


1
Expert's answer
2021-06-02T17:18:00-0400

Solution:

Given, "\mu=300,\sigma=50"

"X\sim N(\mu,\sigma)
\\P(X\ge400)=1-P(X<400)
\\=1-P(z<\dfrac{400-300}{50})
\\=1-P(z<2)
\\=1-0.97725
\\=0.02275"